OpenStudy (anonymous):

If f(0) = 6 and the function f(x) has a property f(2x+3)= 2 f(x) + 3 Find f (3), f(9) and f (-3)

OpenStudy (anonymous):

I believe the properties imply that 2 f(x)= f(2x) and f(x) + 3= f(x+3) This said if we can rewrite those values in terms of x you will be able to find the values at those points if you need more help let me know but I like to try to let people solve the question as much as possible on their own

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

Just remember to medal and close the question.